### 26 Mar 2026 v2

# Neural Uncertainty Principle: A Unified View of Adversarial Fragility and LLM Hallucination

Abstract
Adversarial vulnerability in vision and hallucination in large language models are conventionally viewed as separate problems, each addressed with modality-specific patches. This study first reveals that they share a common geometric origin: the input and its loss gradient are conjugate observables subject to an irreducible uncertainty bound. Formalizing a Neural Uncertainty Principle (NUP) under a loss-induced state, we find that in near-bound regimes, further compression must be accompanied by increased sensitivity dispersion (adversarial fragility), while weak prompt–gradient coupling leaves generation under-constrained (hallucination). Crucially, this bound is modulated by an input–gradient correlation channel, captured by a specifically designed single-backward probe. In vision, masking highly coupled components improves robustness without costly adversarial training; in language, the same prefill-stage probe detects hallucination risk before generating any answer tokens. NUP thus turns two seemingly separate failure taxonomies into a shared uncertainty-budget view and provides a principled lens for reliability analysis. Guided by this NUP theory, we propose ConjMask (masking high-contribution input components) and LogitReg (logit-side regularization) to improve robustness without adversarial training, and use the probe as a decoding-free risk signal for LLMs, enabling hallucination detection and prompt selection. NUP thus provides a unified, practical framework for diagnosing and mitigating boundary anomalies across perception and generation tasks.
I Introduction
Modern neural systems have evolved far beyond their initial roles as pattern recognizers. They have become integral infrastructure in scientific discovery, industrial automation, and daily digital interaction [30, 55, 19, 64, 3, 43, 44, 26]. As these models extend across critical domains, from predicting protein structures to generating legal or medical advice, the requirement to understand their boundary behaviors shifts from a theoretical curiosity to an engineering necessity [67, 23, 22]. This fragility is most clearly manifested in two distinct yet pervasive phenomena: adversarial vulnerability in vision models, where imperceptible perturbations flip confident predictions [61, 16, 33, 38, 4, 2, 7, 6], and hallucination especially in large language models (LLMs), where generation drifts fluently from fact to fabrication [36, 35, 45, 10, 48, 29, 41, 27].
Currently, the dominant responses to these failures have been domain- and modality-specific. Adversarial vulnerability is typically formalized as worst-case risk under norm-bounded perturbations, and the dominant mitigation is adversarial training (AT) and its variants (e.g., TRADES), which improve robustness but are computationally expensive and tightly coupled to specific threat models [33, 68]. Hallucination in LLMs is usually framed as a faithfulness/factuality failure in generation. Mitigation is commonly pursued through alignment and instruction tuning [42], retrieval-augmented generation (RAG) that augments parametric memory with external evidence [32], and post-hoc verification/self-check pipelines [34, 14, 8].
Overall, current practice still largely reflects a patchwork of modality-specific solutions, rather than converging toward a unified observable of boundary behavior that is comparable across perception and generation.
Here, we take a unified view and argue that both failure modes share a geometric signature: the model enters a boundary-stress of the loss landscape. In adversarial attacks, small input perturbations can traverse locally sharp loss contours; in LLMs, weak prompt conditioning leaves a large set of continuations similarly compatible before decoding, enabling prior-driven drift. We show that these are two opposite ways of mismanaging the same “uncertainty budget”—an imbalance between input-space localization and gradient-space sensitivity.
In summary, this paper mainly makes the following contributions:
  • Neural Uncertainty Principle (NUP). We formalize a Neural Uncertainty Principle (NUP) for neural models with a well-defined scalar loss: under a boundary-emphasized (loss-induced) state, an input projection operator and its directional-derivative conjugate obey a Robertson–Schrödinger–type constraint (Theorem III.1). The constraint captures a fundamental trade-off: a model cannot be made simultaneously arbitrarily accurate on boundary-relevant samples and uniformly robust to small input perturbations.
  • From operator constraint to a computable correlation channel. A key technical contribution is that, under the loss-phase construction, the covariance term in the Robertson–Schrödinger (RS) inequality admits an exact reduction to loss-weighted statistics of the real loss gradient field (Theorem III.2, Lemma III.2). This reduction exposes an explicit input–gradient coupling channel that controls proximity to the constraint bound, and yields a practical single-backward per-sample proxy: the Conjugate Correlation Probe (CC-Probe), defined as the absolute input–gradient cosine (mean-centered for prompt embeddings).
  • Two opposite anomaly regimes under the same principle. Using the CC-Probe, we can empirically identify two opposite boundary-anomaly regimes: in vision, a persistent high-CC-Probe marks boundary-stress and concentrates adversarial fragility; in LLM prompting, anomalously low CC-Probe marks under-conditioning at prefill and correlates with elevated hallucination risk. Across settings, reliable behavior typically lies between these extremes, which we summarize as an intermediate “Goldilocks” band (Fig. 1) used as a qualitative regime indicator.
To validate the aforementioned results, we design six experiments that diagnose the two regimes and use two lightweight interventions (ConjMask, LogitReg) as mechanistic tests that manipulate the coupling channel and produce objective-specific shifts under controlled stress tests.
Our goal is to establish NUP as a unified and testable lens for boundary anomalies, by turning its correlation channel into a single-backward observable and validating the resulting regime predictions with controlled mechanistic interventions, rather than simply optimizing an end-to-end adversarial defense or hallucination detector for leaderboards.
II Related Work
II-A Adversarial Robustness and Accuracy–Robustness Frontier
Deep neural networks achieve remarkable accuracy on clean data [30, 19, 64], but remain vulnerable to imperceptible adversarial perturbations [61, 16]. In robustness evaluation, a standard first-order white-box adversary is Projected Gradient Descent (PGD) [33], which iteratively maximizes the training loss—typically the cross-entropy (CE) loss for classification. More recently, standardized suites such as AutoAttack [7] include Auto-PGD (APGD) variants that optimize either CE (APGD-CE) or the Difference of Logits Ratio (DLR) loss (APGD-DLR), offering complementary stress tests beyond plain CE optimization.
Adversarial training (AT) formulates robustness as a min–max optimization problem [33], and TRADES [68] further decomposes the objective to balance natural and robust accuracy. Despite progress, a persistent accuracy–robustness tension remains, attributed variously to sample complexity [50], non-robust features [63], and intrinsic margin constraints [11, 58]. Standardized benchmarks such as RobustBench [6] and strong attacks like AutoAttack [7] now enable rigorous evaluation.
Our work offers a complementary perspective: Proposition 1 frames this tension as a conjugate necessity arising from the NUP. As models compress features to sharpen decision boundaries, the effective feasible volume shrinks, pushing the system toward bound saturation.
II-B Gradient-Based Attribution, Regularization, and Masking
[…]
Masking-based regularization offers another avenue. Dropout [57] and DropBlock [15] randomly mask activations to prevent co-adaptation, while adversarial masking strategies [5] target vulnerable features. Our ConjMask (presented in Sec. V-D) differs fundamentally: rather than random or adversarially-driven selection, we mask input components with large normalized interaction scores $|\tilde{x}_{c,j}\tilde{p}_{c,j}|$ (Eq. (36)), i.e., a channel-wise normalized version of the $x_{i}p_{i}$-type coupling, directly targeting the input–gradient coupling channel that appears in the NUP via the correlation coefficient $\rho_{c}(u)$ (Eq. (15)) and the covariance reduction (Theorem III.2). Meanwhile, prior GxI-based methods use $x\odot\nabla_{x}\mathcal{L}$ for post-hoc explanation; we instead treat $|\cos(x,p)|$ as a predictive state variable motivated by the NUP (Theorem III.1) and its mixed-axis corollary (Lemma III.1).
II-C Uncertainty Quantification and Conjugate Perspectives
Predictive uncertainty estimation enables models to flag unreliable outputs. Deep Ensembles [31] remain a strong baseline, while Monte Carlo Dropout [12] provides a Bayesian approximation. Calibration techniques [18] align confidence with accuracy, critical for out-of-distribution detection [21, 67]. These methods typically require multiple forward passes.
A separate line explores trade-off structures echoing conjugate relationships. The Information Bottleneck [62] formalizes compression–prediction trade-offs in representation space; Spectral Normalization [37] and gradient penalties [17] implicitly constrain sensitivity dispersion. Logit regularization and label smoothing [60, 39] stabilize output distributions against overconfidence.
Our NUP framework differs in grounding uncertainty in input-space conjugate geometry rather than representation-level compression or output-level statistics. The Robertson–Schrödinger form (Theorem III.1) yields a single-backward probe that complements expensive sampling-based estimators.
II-D Hallucination Detection and Mitigation in LLMs
Large language models generate fluent but factually unsupported text—a phenomenon termed hallucination [36, 35, 48]. Recent theoretical analyses suggest hallucinations may be structurally unavoidable under realistic conditions [41, 27].
Detection approaches are predominantly post-hoc and sampling-based. Semantic entropy methods [10, 29] estimate uncertainty via multiple generations, while LLM-as-a-judge [72] uses auxiliary models to evaluate outputs. Prompt engineering strategies such as Chain-of-Thought [66, 28] reshape the generation process itself. Self-consistency decoding [65] aggregates multiple reasoning paths but requires substantial sampling overhead.
These methods share a common limitation: they operate during or after decoding. Early-exit mechanisms [52] and confidence-based filtering [25] reduce inference cost but still require partial generation. In contrast, our prefill-stage probe (presented in Sec. V-FV-G) computes $|\cos(\bar{x},\bar{p})|$ on prompt embeddings before generating any answer (output) tokens, enabling decoding-free and sampling-free risk assessment. Proposition 4 interprets this through the NUP: anomalously low correlation indicates a high-slack regime (i.e., $S_{c}(u)\gg 1$ in Eq. (28)), where the prompt fails to constrain the feasible continuation space, amplifying prior-driven drift.
II-E Positioning of This Work
Existing approaches address adversarial robustness and hallucination as separate phenomena, employing modality-specific defenses (adversarial training, prompt engineering) or post-hoc diagnostics (semantic entropy, LLM-as-judge). Our work departs from this paradigm in the following manners:
  1. (i)
    Unified mechanism. We ground both failure modes in a single operator-theoretic constraint (Theorem III.1), where the input–gradient correlation $\rho_{c}(u)$ governs proximity to the uncertainty bound.
  2. (ii)
    Prefill signal. Unlike sampling-based uncertainty estimators, the CC-Probe requires only a single backward pass and operates before decoding (for LLMs) or attack evaluation (for vision).
  3. (iii)
    Intervention design. Rather than adversarially augmenting the training set, we directly target the conjugate coupling channel that the NUP identifies as the source of fragility.
III Formalism of NUP
III-A NUP in a Nutshell: Roadmap From Geometry to Experiments
Figure 1: Illustration of the Neural Uncertainty Principle (NUP) in the $(\Delta\hat{m}_{u}^{\star},\Delta\hat{p}_{u})$ plane. Here $\Delta\hat{m}_{u}^{\star}$ is the minimum dispersion achievable by a mixed observable $\hat{m}_{u}(\lambda)=\hat{x}_{u}+\lambda\hat{p}_{u}$, and $\Delta\hat{p}_{u}$ is the dispersion of the conjugate operator $\hat{p}_{u}$ under the loss-phase (boundary-emphasized) state. NUP implies the forbidden region $\Delta\hat{m}_{u}^{\star}\Delta\hat{p}_{u}<\tfrac{1}{2}$. Vision boundary-stress concentrates toward the upper-left (small $\Delta\hat{m}_{u}^{\star}$, large $\Delta\hat{p}_{u}$), whereas LLM under-conditioning / hallucination occupies a larger region away from the uncertainty boundary (large $\Delta\hat{m}_{u}^{\star}$, uncontrolled $\Delta\hat{p}_{u}$). Reliable behavior lies in an intermediate “Goldilocks” band where neither term is extreme. Note. Strictly speaking, $\Delta\hat{m}_{u}^{\star}$ and $\Delta\hat{p}_{u}$ denote the dispersions of the operators $\hat{m}_{u}$ and $\hat{p}_{u}$ under the loss-induced state (hence the “hat” notation at the operator level). In the figure, to emphasize the phase-plane geometry and the experimentally observable quantities, we plot numerical estimates of these dispersions. We therefore omit the hats in the figure labels for notational simplicity; the plotted quantities correspond one-to-one to $\Delta\hat{m}_{u}^{\star}$ and $\Delta\hat{p}_{u}$ as defined in the equations.
[…]
  • $\Delta\hat{p}_{u}$ (sensitivity dispersion): an operator dispersion whose real drift term is driven by the loss gradient. Operationally, it quantifies how widely the boundary-emphasized population spreads in loss-sensitivity: larger $\Delta\hat{p}_{u}$ indicates that small input perturbations can more frequently induce large loss changes, which is precisely the regime exploited by gradient-based attacks.
  • $\Delta\hat{m}_{u}^{\star}$ (best mixed-axis thickness): define the mixed observable $\hat{m}_{u}(\lambda)=\hat{x}_{u}+\lambda\hat{p}_{u}$ and let $\Delta\hat{m}_{u}(\lambda)$ be its dispersion under the same state. Then $\Delta\hat{m}_{u}^{\star}:=\min_{\lambda}\Delta\hat{m}_{u}(\lambda)$ is the minimum achievable thickness of the boundary-emphasized ellipse along any mixed (tilted) axis. Intuitively, it measures how consistent the boundary layer is across samples: a smaller $\Delta\hat{m}_{u}^{\star}$ means the hard population collapses into a thinner ambiguity band in the phase plane, which typically accompanies higher clean accuracy but also sharper boundary behavior.
Mixed-axis NUP constraint. The full NUP is a Robertson–Schrödinger uncertainty relation for the operator pair $(\hat{x}_{u},\hat{p}_{u})$ (Theorem III.1). A key corollary (Lemma III.1) eliminates the covariance term by passing to the optimal mixed axis and yields
$\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\;\geq\;\tfrac{1}{2}.$ (1)
This inequality is the only constraint required to interpret Fig. 1: the lower-left region is forbidden, i.e., one cannot simultaneously make the boundary-emphasized population extremely thin (very small $\Delta\hat{m}_{u}^{\star}$) and uniformly insensitive (very small $\Delta\hat{p}_{u}$). This is the geometric origin of adversarial vulnerability and hallucination in view of the NUP.
Two opposite failure modes on the same plane. The same $(\Delta\hat{m}_{u}^{\star},\Delta\hat{p}_{u})$ plane explains two “boundary anomalies” as opposite corners:
Vision boundary-stress
Discriminative training tends to squeeze errors and near-margin samples into a narrow boundary layer, which corresponds to pushing $\Delta\hat{m}_{u}^{\star}$ downward. Based on Eq. (1), we know that this compression must be compensated by inflating $\Delta\hat{p}_{u}$: the hard subset becomes increasingly sensitive. Empirically, this is where wrong / semi-hard samples concentrate—they are highly correctable by small perturbations, hence adversarially fragile.
LLM under-conditioning / hallucination
In prompting, a common failure is not an overly sharp boundary but an under-constrained continuation space. If the prompt does not sufficiently couple to loss-sensitive directions at prefill, the effective phase-plane cloud is thick (large $\Delta\hat{m}_{u}^{\star}$) while sensitivity dispersion is unbounded (uncontrolled $\Delta\hat{p}_{u}$), corresponding to high slack: many continuations are similarly compatible before decoding begins. This “under-conditioning” corner aligns with elevated hallucination risk, where generation can drift into fluent but weakly grounded outputs.
Intermediate “Goldilocks” zone
Reliable behavior lies between extremes: $\Delta\hat{m}_{u}^{\star}$ is not excessively squeezed (avoiding boundary-stress), and $\Delta\hat{p}_{u}$ is not vanishingly small under weak conditioning (avoiding unconstrained drift). We will repeatedly interpret experiments as moving populations toward this intermediate band in Fig. 1.
Practical probe. The Robertson–Schrödinger form contains an explicit correlation term $\rho_{c}(u)$ and satisfies $\Delta\hat{m}_{u}^{\star}=\Delta\hat{x}_{u}\sqrt{1-\rho_{c}(u)^{2}}$ (Eq. 20), where $\rho$ controls the tilt/coupling of the ellipse, while the primary “budget” relevant to failure modes is captured by $(\Delta\hat{m}_{u}^{\star},\Delta\hat{p}_{u})$ through Eq. (1).
To localize individual samples on this plane without estimating operator moments, we use a single-backward proxy for the coupling channel: the input–gradient cosine (CC-Probe), i.e., $c_{\text{img}}$ in vision and $c_{\text{prompt}}$ in LLM prefill (Eqs. (33)–(34)). It tracks the same input–gradient coupling that enters the Robertson–Schrödinger covariance term under the loss-phase construction, and empirically identifies boundary-stress in vision and under-conditioning prompts in language.
Preview our experiment objectives using Fig. 1.
[…]
  • Exp. 3–4 (vision intervention): ConjMask suppresses dominant $|x_{i}p_{i}|$ couplings during training (relieving boundary stress); LogitReg complements it against score-space attacks. Net effect: shift the stressed population away from the upper-left corner.
  • Exp. 5–6 (LLM intervention): low $c_{\text{prompt}}$ indicates under-conditioning at prefill; selecting higher-$c_{\text{prompt}}$ prompt variants moves behavior towards the Goldilocks Zone.
In short, NUP forbids the lower-left: one cannot simultaneously minimize boundary ambiguity and sensitivity. Vision failures arise when training pushes the hard subset toward small $\Delta\hat{m}_{u}^{\star}$ (hence large $\Delta\hat{p}_{u}$), whereas LLM hallucinations arise under high slack with large $\Delta\hat{m}_{u}^{\star}$ and uncontrolled $\Delta\hat{p}_{u}$. Our probes and interventions are designed to steer behavior toward the intermediate band in Fig. 1.
A note to the reader: The following subsections present an operator/commutator formulation (borrowed from the Robertson–Schrödinger uncertainty relation) to justify the correlation channel behind our probe. Readers could safely skip ahead to Secs. IVV for the experimental results and practical algorithms, and return to this theory section for context as needed.
III-B Loss-weighted state and canonical directional operators
Let $f_{\theta}$ be a differentiable model and $\mathcal{L}(\cdot)$ a scalar loss. For a condition $c$, define the per-sample loss
$\mathcal{L}_{c}(x):=\mathcal{L}\big(f_{\theta}(x),c\big).$ (2)
The real input-gradient field (loss sensitivity) is
$p(x):=\nabla_{x}\mathcal{L}_{c}(x)\in\mathbb{R}^{d}.$ (3)
For small $\delta$, $\mathcal{L}_{c}(x+\delta)-\mathcal{L}_{c}(x)=p(x)^{\top}\delta+o(\|\delta\|_{2})$, so $p(x)$ is the standard first-order sensitivity object.
Loss-induced weighting and state. To emphasize boundary-relevant regions in operator moments, we evaluate second-order statistics under a loss-induced analysis weighting
$w_{c}(x):=\frac{\mathcal{L}_{c}(x)^{2}}{\int_{\mathcal{X}}\mathcal{L}_{c}(\xi)^{2}\,d\xi},\qquad\mathbb{E}_{c}[g]:=\int_{\mathcal{X}}g(x)\,w_{c}(x)\,dx.$ (4)
In practice, all expectations are instantiated either over an empirical dataset distribution or over the prompt set used in evaluation.
It is convenient to encode $w_{c}$ as a normalized state in $\mathcal{H}:=L^{2}(\mathcal{X})$:
$\psi_{c}(x)$= $A_{c}(x)\exp\!\big(i\alpha\,\mathcal{L}_{c}(x)\big),$ (5)
$A_{c}(x)$= $\frac{|\mathcal{L}_{c}(x)|}{\sqrt{\beta_{c}}},$
$\beta_{c}$= $\int_{\mathcal{X}}\mathcal{L}_{c}(x)^{2}\,dx,$
so that $|\psi_{c}(x)|^{2}=A_{c}(x)^{2}=w_{c}(x)$. We fix the phase scale $\alpha=1$ throughout. It only sets the phase convention of the auxiliary state and is not treated as a tunable parameter. For a (densely defined) operator $\hat{O}$ on $L^{2}(\mathcal{X})$, we write its expectation [49] as  [^2] The expectation $\langle\hat{O}\rangle_{c}$ is the average value of this quantity in the state $\psi_{c}$, computed via the usual $L^{2}$ inner product.
$\langle\hat{O}\rangle_{c}\;:=\;\langle\psi_{c},\hat{O}\psi_{c}\rangle\;=\;\int_{\mathcal{X}}\psi_{c}(x)^{*}\,(\hat{O}\psi_{c})(x)\,dx,$ (6)
where $(\cdot)^{*}$ denotes complex conjugation.
Directional projection. To avoid ambiguity in vector-valued uncertainty statements, we introduce a unit direction $u\in\mathbb{S}^{d-1}$ and study one-dimensional projections:
$x_{u}$= $u^{\top}x,$ (7)
$p_{u}(x)$= $u^{\top}p(x)=\partial_{u}\mathcal{L}_{c}(x),\qquad\partial_{u}=u^{\top}\nabla_{x}.$
Canonical operators. Define the multiplication and directional derivative operators as follows:
$(\hat{x}_{u}g)(x):=(u^{\top}x)\,g(x),\qquad(\hat{p}_{u}g)(x):=-i\,\partial_{u}g(x),$ (8)
on a common dense domain $\mathcal{D}\subset L^{2}(\mathcal{X})$ where the following algebraic manipulations are valid. A direct calculation on $\mathcal{D}$ gives, for any test function $g$,
$[\hat{x}_{u},\hat{p}_{u}]g$ $=\hat{x}_{u}(-i\partial_{u}g)-(-i\partial_{u})((u^{\top}x)g)$
$=-i(u^{\top}x)\partial_{u}g+i\big((u^{\top}x)\partial_{u}g+g\big)=ig,$ (9)
hence
$\hat{x}_{u},\hat{p}_{u}]=i\mathbb{I}.$ (10)
We use the canonical constant
$\kappa:=\tfrac{1}{2}\big|\langle[\hat{x}_{u},\hat{p}_{u}]\rangle_{c}\big|=\tfrac{1}{2},$ (11)
which follows immediately from (10) and $\langle\mathbb{I}\rangle_{c}=1$. Eq. (10) is similar to the canonical position–momentum commutation relation from quantum mechanics [20, 9].
III-C RS form of NUP and the mixed-axis corollary
Operator inequalities constrain fluctuations rather than means. Accordingly, for an observable $\hat{O}$ we write its centered version as $\Delta\hat{O}:=\hat{O}-\langle\hat{O}\rangle_{c}$. The dispersions of $\hat{x}_{u}$ and $\hat{p}_{u}$ under $\psi_{c}$ are then
$(\Delta\hat{x}_{u})^{2}$= $\langle(\Delta\hat{x}_{u})^{2}\rangle_{c}=\mathbb{E}_{c}\!\big[(x_{u}-\mathbb{E}_{c}[x_{u}])^{2}\big],$ (12)
$(\Delta\hat{p}_{u})^{2}$= $\langle(\Delta\hat{p}_{u})^{2}\rangle_{c}=\big\|\Delta\hat{p}_{u}\,\psi_{c}\big\|_{2}^{2},$
and the symmetrized covariance is
$\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u}):=\frac{1}{2}\Big\langle\Delta\hat{x}_{u}\,\Delta\hat{p}_{u}+\Delta\hat{p}_{u}\,\Delta\hat{x}_{u}\Big\rangle_{c}.$ (13)
Theorem III.1 (Neural Uncertainty Relation).
For the canonical pair $(\hat{x}_{u},\hat{p}_{u})$ and the state $\psi_{c}$ with finite second moments,
$(\Delta\hat{x}_{u})^{2}(\Delta\hat{p}_{u})^{2}\ \geq\ \kappa^{2}+\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})^{2},$ (14)
where $\kappa=\tfrac{1}{2}\big|\langle[\hat{x}_{u},\hat{p}_{u}]\rangle_{c}\big|=\tfrac{1}{2}$.
The inequality is the classical Robertson–Schrödinger relation in quantum physics [46, 51]; a standard proof is provided in the Supplementary Material.
Correlation channel. Define the (operator) correlation coefficient
$\rho_{c}(u):=\frac{\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})}{\Delta\hat{x}_{u}\,\Delta\hat{p}_{u}}\in[-1,1].$ (15)
Then Theorem III.1 is equivalently
$\big(\Delta\hat{x}_{u}\,\Delta\hat{p}_{u}\big)^{2}\bigl(1-\rho_{c}(u)^{2}\bigr)\ \geq\ \kappa^{2}.$ (16)
The factor $(1-\rho^{2})$ is the explicit coupling term that will later be approximated by a cheap proxy.
Mixed-axis elimination of covariance. A useful way to “spend” the covariance term is to pass to a mixed observable. For $\lambda\in\mathbb{R}$, define
$\hat{m}_{u}(\lambda):=\hat{x}_{u}+\lambda\,\hat{p}_{u}.$ (17)
Expanding the variance yields
$(\Delta\hat{m}_{u}(\lambda))^{2}=(\Delta\hat{x}_{u})^{2}+\lambda^{2}(\Delta\hat{p}_{u})^{2}+2\lambda\,\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u}).$ (18)
The minimizing coefficient is
$\lambda^{\star}=-\frac{\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})}{(\Delta\hat{p}_{u})^{2}},$ (19)
and the minimum mixed variance satisfies the identity
$(\Delta\hat{m}_{u}^{\star})^{2}:=\min_{\lambda}(\Delta\hat{m}_{u}(\lambda))^{2}=(\Delta\hat{x}_{u})^{2}\bigl(1-\rho_{c}(u)^{2}\bigr).$ (20)
Lemma III.1 (Mixed-axis form of NUP).
With $\Delta\hat{m}_{u}^{\star}$ defined by (20), Theorem III.1 implies
$\boxed{\ \Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\ \geq\ \kappa=\tfrac{1}{2}\ }.$ (21)
Proof. Substitute (20) into (16) to obtain $(\Delta\hat{m}_{u}^{\star})^{2}(\Delta\hat{p}_{u})^{2}\geq\kappa^{2}$, and take square roots. $\square$
When $\rho_{c}(u)=0$, (20) gives $\Delta\hat{m}_{u}^{\star}=\Delta\hat{x}_{u}$ and Lemma III.1 reduces to the canonical Heisenberg form
$\Delta\hat{x}_{u}\,\Delta\hat{p}_{u}\ \geq\ \tfrac{1}{2}.$ (22)
In this regime the mixed axis offers no thinner direction than $\hat{x}_{u}$ itself; this is the coupling-free case aligned with the Heisenberg-type tension studied in [71, 24, 69, 70].
III-D Interpreting Lemma III.1: ellipse geometry and conjugate trade-off
It is convenient to switch between (i) the operator notation $(\hat{x}_{u},\hat{p}_{u})$ defining second moments under the state $\psi_{c}$, and (ii) the induced scalar random variables obtained by evaluating these observables on samples. Concretely, $\hat{x}_{u}$ is a multiplication operator, so its “measurement” on a sample $x$ is the scalar $x_{u}=u^{\top}x$; accordingly, $\langle\hat{x}_{u}\rangle_{c}=\mathbb{E}_{c}[x_{u}]$ and $(\Delta\hat{x}_{u})^{2}=\mathbb{E}_{c}[(x_{u}-\mathbb{E}_{c}[x_{u}])^{2}]$. For $\hat{p}_{u}=-i\partial_{u}$, the corresponding real sensitivity channel is the directional loss gradient $p_{u}(x)=\partial_{u}\mathcal{L}_{c}(x)$. Under the loss-induced state, the symmetrized operator covariance reduces exactly to a scalar covariance (Theorem III.2).
Boundary layer induced by loss weighting. All expectations $\mathbb{E}_{c}[\cdot]$ are taken under the analysis weighting $w_{c}(x)\propto\mathcal{L}_{c}(x)^{2}$ (Section III-B). In classification, large loss concentrates on misclassified or small-margin samples, which are typically boundary-adjacent. Therefore, the loss-weighted population emphasized by $w_{c}$ can be viewed as an effective boundary layer.
Ellipse view of the loss-weighted $(x_{u},p_{u})$ cloud. With the above identification, the loss-weighted samples form a 2D cloud of points $(x_{u},p_{u})$ in the $(x_{u},p_{u})$ plane. Its second-order geometry is summarized by the covariance matrix
$\Sigma(u):=\begin{pmatrix}(\Delta\hat{x}_{u})^{2}&\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})\\ \mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})&(\Delta\hat{p}_{u})^{2}\end{pmatrix}$. (23)
Geometrically, $\Sigma(u)$ defines an ellipse: the correlation $\rho_{c}(u)$ controls its tilt (coupling), while $\det\Sigma(u)$ controls its area. A short calculation gives
$\det\Sigma(u)=$ $(\Delta\hat{x}_{u})^{2}(\Delta\hat{p}_{u})^{2}-\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})^{2}$ (24)
$=$ $\big(\Delta\hat{x}_{u}\,\Delta\hat{p}_{u}\big)^{2}\bigl(1-\rho_{c}(u)^{2}\bigr).$
Combining (24) with (20) yields the factorization
$\sqrt{\det\Sigma(u)}=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u},$ (25)
so Theorem III.1 and Lemma III.1 can be read as an area lower bound:
$\sqrt{\det\Sigma(u)}\ \geq\ \kappa=\tfrac{1}{2}\ \Longleftrightarrow\ \Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\ \geq\ \tfrac{1}{2}.$ (26)
Stability of the loss landscape. The mixed observable $\hat{m}_{u}(\lambda)=\hat{x}_{u}+\lambda\hat{p}_{u}$ corresponds to projecting the $(x_{u},p_{u})$ cloud onto the line $m_{u}(\lambda)=x_{u}+\lambda p_{u}$. The quantity $\Delta\hat{m}_{u}^{\star}:=\min_{\lambda}\Delta\hat{m}_{u}(\lambda)$ is the smallest standard deviation achievable among this one-parameter family of projections, i.e., the thickness of the loss-weighted cloud along its sharpest mixed axis.
Thus, the decrease of $\Delta\hat{m}_{u}^{\star}$ means that within the loss-weighted (boundary-layer) population, the position–sensitivity relation becomes more consistent: similar positions $x_{u}$ tend to induce similar sensitivities $p_{u}$, and conversely, a given sensitivity level $p_{u}$ corresponds to a narrower range of positions $x_{u}$. Equivalently, the local loss landscape seen by boundary-adjacent samples is more “aligned” across samples, producing a thinner ambiguity band in the $(x_{u},p_{u})$ phase plane. We emphasize that this is an effective thickness (a second-order spread of a loss-weighted phase-plane population), not the Euclidean thickness of the decision boundary in input space.
Connection to clean accuracy. Since natural errors concentrate near the decision boundary, and $w_{c}$ upweights precisely these boundary-adjacent/high-loss samples, reducing this boundary-layer ambiguity band (smaller $\Delta\hat{m}_{u}^{\star}$) typically reduces near-boundary natural misclassifications. Consequently, $\Delta\hat{m}_{u}^{\star}$ often co-varies inversely with clean accuracy along training trajectories.
Conjugate axis as sensitivity dispersion and robustness. In contrast, $\Delta\hat{p}_{u}$ measures the dispersion of the derivative operator $\hat{p}_{u}$ under $\psi_{c}$; under the loss-induced state, $\hat{p}_{u}\psi_{c}$ contains a real drift term proportional to the real gradient $p_{u}(x)$ (Section III-E), so $\Delta\hat{p}_{u}$ captures a boundary-weighted sensitivity scale. This connects to robustness via the first-order expansion
$\mathcal{L}_{c}(x+\delta)-\mathcal{L}_{c}(x)=p(x)^{\top}\delta+o(\|\delta\|_{2}),$ (27)
which implies that, under an $L_{2}$-bounded perturbation $\|\delta\|_{2}\leq\epsilon$, the worst-case first-order loss increase scales as $\epsilon\|p(x)\|_{2}$. Therefore, larger boundary-weighted sensitivity dispersion (i.e., larger $\Delta\hat{p}_{u}$) indicates that smaller perturbations tend to induce larger loss changes and more frequent boundary crossings, degrading adversarial robustness under gradient-based threat models.
Accuracy–robustness tension as a conjugate constraint. Eq. (26) is the central geometric message: when the ellipse area is close to its RS lower bound, reducing the thin-axis thickness $\Delta\hat{m}_{u}^{\star}$ (sharper boundary-layer consistency / thinner ambiguity band) necessarily enlarges the conjugate sensitivity dispersion $\Delta\hat{p}_{u}$. This is the sense in which the two quantities are conjugate: compressing the boundary-weighted $(x_{u},p_{u})$ cloud along its sharpest mixed axis must be compensated by expansion along the sensitivity axis so that the area constraint remains satisfied. This provides a principled geometric vocabulary for the empirical accuracy–robustness tension.
It is sometimes convenient to define
$V_{c}(u):=\sqrt{\det\Sigma(u)}=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u},\ \ S_{c}(u):=\frac{V_{c}(u)}{\kappa}\geq 1.$ (28)
Here $S_{c}(u)\approx 1$ indicates a near-bound regime (limited slack), where the conjugate trade-off is most pronounced, while $S_{c}(u)\gg 1$ indicates substantial slack in the second-order geometry.
III-E From operator moments to real gradients and a practical proxy
The NUP above is stated at the operator level. To connect it to computable neural quantities, we first show that the operator covariance $\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})$ reduces exactly to a scalar covariance involving the real gradient $p_{u}(x)$ under the loss-phase state (5). We then introduce a cheap proxy for the coupling channel.
Operator covariance reduction
We first present the following theorem.
Theorem III.2 (Covariance reduction under the loss-induced state).
Let $\psi_{c}(x)=A_{c}(x)e^{i\alpha\mathcal{L}_{c}(x)}$ with $A_{c}^{2}=w_{c}$ as defined in Eq. (5). Assume that boundary terms vanish for integration by parts along direction $u$, and then the symmetrized operator covariance satisfies
$\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})=\alpha\,\mathrm{Cov}^{\mathrm{sc}}_{c}(x_{u},p_{u}),$ (29)
where $\mathrm{Cov}^{\mathrm{sc}}_{c}(g,h):=\mathbb{E}_{c}[gh]-\mathbb{E}_{c}[g]\mathbb{E}_{c}[h]$ is the scalar covariance under $w_{c}$ and $p_{u}(x)=\partial_{u}\mathcal{L}_{c}(x)$.
Proof sketch. We rewrite the argument in terms of the centered variable $\Delta x_{u}$ and apply the drift identity for $i\partial_{u}\psi_{c}$, which makes the $\partial_{u}A_{c}$ contribution purely imaginary and hence irrelevant after taking the real part in the symmetrized covariance. Please see more proof details in the supplementary material.
CC-Probe as a proxy for the coupling channel
The operator correlation $\rho_{c}(u)$ is defined through dataset-level moments under $w_{c}$ and depends on the choice of direction $u$. To obtain a cheap per-sample statistic that tracks the same coupling channel appearing in $\mathrm{Cov}^{\mathrm{sc}}_{c}(x_{u},p_{u})$, we use a geometric identity for random directional probes.
Lemma III.2 (Directional-probe correlation and cosine).
Fix nonzero vectors $x,p\in\mathbb{R}^{d}$, and draw $u\sim\mathrm{Unif}(\mathbb{S}^{d-1})$. Let $x_{u}:=u^{\top}x$ and $p_{u}:=u^{\top}p$. Then
$\mathrm{Corr}_{u}(x_{u},p_{u})=\frac{x^{\top}p}{\|x\|_{2}\,\|p\|_{2}}=\cos(x,p).$ (30)
Proof. Since $\mathbb{E}_{u}[u]=0$ and $\mathbb{E}_{u}[uu^{\top}]=\frac{1}{d}I$,
$\mathbb{E}_{u}[x_{u}p_{u}]=$ $\mathbb{E}_{u}[(u^{\top}x)(u^{\top}p)]=x^{\top}\mathbb{E}_{u}[uu^{\top}]p=\tfrac{1}{d}x^{\top}p,$ (31)
$\mathbb{E}_{u}[x_{u}^{2}]=$ $\tfrac{1}{d}\|x\|_{2}^{2},$
$\mathbb{E}_{u}[p_{u}^{2}]=$ $\tfrac{1}{d}\|p\|_{2}^{2}.$
Because $\mathbb{E}_{u}[x_{u}]=\mathbb{E}_{u}[p_{u}]=0$, the Pearson correlation equals $\mathrm{Corr}_{u}(x_{u},p_{u})=\mathbb{E}_{u}[x_{u}p_{u}]/\sqrt{\mathbb{E}_{u}[x_{u}^{2}]\mathbb{E}_{u}[p_{u}^{2}]}=\cos(x,p)$. $\square$
Motivated by Lemma III.2 and Theorem III.2, we provide the Conjugate Correlation Probe (CC-Probe, briefly) for a sample $x$ as
$c_{\text{probe}}(x):=\big|\cos\!\big(x,\,p(x)\big)\big|=\frac{\big|x^{\top}p(x)\big|}{\|x\|_{2}\,\|p(x)\|_{2}}.$ (32)
The quantity $c_{\text{probe}}$ is a cheap per-sample indicator of directional coupling: averaged over random directions, $\cos(x,p)$ is exactly the correlation between the projected coordinate and projected gradient (Lemma III.2), and the dataset-level covariance channel driving $\rho_{c}(u)$ reduces to scalar covariances between $x_{u}$ and $p_{u}$ (Theorem III.2). This relationship is sufficient for $c_{\text{probe}}$ to serve as a practical proxy used in our experiments and interventions.
III-F Testable predictions
The NUP motivates a simple idea: abnormal input–gradient coupling is a reliable marker of failure modes. We test this idea with the CC-Probe introduced above.
Proposition 1 (Exp. 1 prediction: a “high-cosine tail” marks hard/fragile vision samples).
As training improves clean accuracy, correctly classified images should show lower CC-Probe $c_{\text{img}}$, while misclassified/hard images remain at higher $c_{\text{img}}$ (a persistent high-cosine tail). Adversarial errors should concentrate on this high-cosine tail.
[…]
For a fixed vision model, a small gradient-aligned perturbation (+FGSM) should increase $c_{\text{img}}$ and reduce accuracy, while a sufficiently small anti-aligned perturbation (–FGSM) should decrease $c_{\text{img}}$ and can preserve (or slightly improve) accuracy.
Proposition 3 (Exp. 3–4 prediction: training that suppresses strong $x\!\cdot\!p$ coupling becomes more robust).
If vulnerability is driven by a few input components with large coupling scores (implemented in practice as the channel-wise normalized interaction $|\tilde{x}_{c,j}\tilde{p}_{c,j}|$), then masking these dominant components during training (ConjMask) should improve robustness to standard gradient attacks (e.g., PGD/APGD-CE) without adversarial training (Exp. 3). If this robustness is loss-dependent, adding a logit stabilizer (LogitReg) should restore robustness under stronger loss-optimized attacks (e.g., APGD-DLR) (Exp. 4).
Proposition 4 (Exp. 5–6 prediction: in LLM prefill, low CC-Probe means under-conditioning and higher hallucination risk).
Using only prefill (no decoding), prompts with unusually low $c_{\text{prompt}}$ should have higher hallucination risk; therefore $c_{\text{prompt}}$ should predict hallucination above chance (Exp. 5). Among paraphrased prompts, choosing the one with higher $c_{\text{prompt}}$ should more often select the judge-preferred prompt (Exp. 6).
IV Experimental Protocol
The preceding section derived four testable predictions from the NUP. We design Experiments 1–6 to test these predictions across two modalities: discriminative vision (classification) and generative language (reasoning). Detailed training recipes and hyperparameters are deferred to Supplementary Material.
IV-A Datasets and Model Zoo
Vision
We cover three levels of visual complexity: (i) CIFAR-10 ($32{\times}32$, coarse-grained); (ii) Tiny-ImageNet-200 ($64{\times}64$, medium-scale); and (iii) ImageNet-100 ($224{\times}224$, a class-balanced subset of ILSVRC2012). Across these benchmarks, we instantiate a diverse “architecture zoo” to ensure architecture-agnostic validity:
  • CNNs: ResNet-18/50, DenseNet-121, EfficientNet-B0/B4.
  • Transformers: Vision Transformer ViT-Tiny and Swin-Tiny.
  • State-Space Models: Vision Mamba (Vim-Tiny) [73].
Language
We employ a publicly available model, deepseek-coder-7b-instruct-v1.5, as the subject model. The evaluation focuses on undergraduate-level math reasoning tasks where hallucination is prevalent. We curated two specific datasets:
  • Benchmark-500: A collection of 500 unique, undergraduate-level mathematics problems.
  • Perturbation-100: A focused subset of 100 problems sampled from Benchmark-500. For each problem, we generated 4 additional semantically equivalent but syntactically distinct rephrasings (5 variations total per problem).
These datasets are designed for controlled falsification of Proposition 4: math reasoning reduces confounds from retrieval and enables more stable verification.
Ground truth for these generative tasks is established via a “Panel-of-Judges” consensus mechanism involving five external LLMs (Claude-4.5-Opus-Think, DeepSeek-v3.2-Think, Gemini-3-Pro, GPT-5.2-Thinking, and Grok-4.1-Think).
IV-B Operationalizing the Conjugate Probe
A core contribution of this work is translating the theoretical operator correlation $\rho_{c}$ into a computable per-sample observable.
Vision
For an image input $X\in\mathbb{R}^{C\times H\times W}$, we compute the gradient of the task loss $\mathcal{L}$ with respect to the input, $P=\nabla_{X}\mathcal{L}$. Both $X$ and $P$ are flattened into 1-D vectors $x$ and $p$. Because the practical vision probe is computed on the normalized input tensor fed to the model, we use the uncentered image-space cosine; the sample-level conjugate correlation and its dataset-level average over $N$ samples are defined as:
$c_{\text{img}}=\frac{|x^{\top}p|}{\|x\|_{2}\|p\|_{2}},\qquad\bar{c}_{\text{img}}=\frac{1}{N}\sum_{i=1}^{N}c_{\text{img}}^{(i)},$ (33)
where $c_{\text{img}}^{(i)}$ is calculated on the $i$-th sample. In practice, unless explicitly stated, we report simple (uniform) averages over a given subset (e.g., correct vs. incorrect samples) rather than explicitly reweighting by $\mathcal{L}^{2}$. The loss-induced weighting is the theoretical device used to derive the operator statistics; empirically, the per-sample cosine serves as a lightweight proxy for input–gradient coupling.
Remark. The NUP is expressed via an operator correlation $\rho_{c}(u)$ under a loss-induced state, while we use the practical proxy $c_{\text{img}}$. $c_{\text{img}}$ is intended to track the input–gradient coupling channel that enters the Robertson–Schrödinger covariance term: under the loss-phase construction, $\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})$ reduces exactly to a (loss-weighted) scalar covariance between $x_{u}$ and $p_{u}$ (Theorem III.2), and the dot-product $x^{\top}p$ is the direction-averaged form of this coupling.
Language
For language models, we extract the signal during the prompt processing stage (prefill). We compute a mean-centered cosine between the prompt embeddings and their input gradients. Let $X\in\mathbb{R}^{T\times d}$ be the embedding tensor for the $T$ valid tokens, and $P\in\mathbb{R}^{T\times d}$ be the corresponding gradient tensor. We use a shifted loss, i.e., the next-token negative log-likelihood. Unless otherwise stated, the shifted NLL loss is computed on the non-system prefill tokens of the chat template (i.e., the user prompt together with template tokens available before answer generation), so the score is available at prefill time without generating any answer tokens; it still requires one backward pass with respect to the prompt embeddings. We flatten these tensors into 1-D vectors $x,p\in\mathbb{R}^{Td}$ and define the mean-centered vectors
$\bar{x}:=x-\mu_{x}\mathbf{1},\qquad\bar{p}:=p-\mu_{p}\mathbf{1},$
where $\mu_{x}=\mathrm{mean}(x)$ and $\mu_{p}=\mathrm{mean}(p)$ are the corresponding scalar means. The prompt-level conjugate score and its corpus-level average over $M$ prompts are computed as:
$c_{\text{prompt}}=\frac{|\bar{x}^{\top}\bar{p}|}{\|\bar{x}\|_{2}\|\bar{p}\|_{2}},\qquad\bar{c}_{\text{prompt}}=\frac{1}{M}\sum_{j=1}^{M}c_{\text{prompt}}^{(j)}.$ (34)
where $c_{\text{prompt}}^{(j)}$ is calculated on the $j$-th prompt.
IV-C Evaluation Metrics
Adversarial Robustness (Vision)
To assess robustness under complementary white-box threat models, we evaluate three widely used $L_{\infty}$ attacks at the target budget: (i) Iterative baseline attack: PGD-20 as the standard multi-step first-order attack. (ii) AutoAttack component (CE loss): APGD-CE (Auto-PGD optimizing cross-entropy) to provide a stronger, well-tuned first-order baseline. (iii) AutoAttack component (DLR loss): APGD-DLR (Auto-PGD optimizing the Difference of Logits Ratio loss) to stress-test robustness beyond CE-optimized gradients.
Hallucination Analysis (Language)
We treat Exp. 5 hallucination detection as a binary classification task (($1=$ hallucination, $0=$ non-hallucination)), quantified by the Area Under the ROC Curve (AUROC) with bootstrap resampling ($N_{\text{boot}}=2000$). For Exp. 6 prefill-stage prompt selection, the primary metrics are Top-1 Hit Rate (frequency with which the selected prompt belongs to the judge-best set) and Mean Regret (the average quality gap between the selected prompt and the best judge score within the same five-prompt cluster). Deterministic decoding is used for reproducibility.
V Experiments and Evidence
V-A Experimental Roadmap
Our experiments test the NUP predictions using a simple, computable proxy—the CC-Probe—across vision classification and LLM prompting. We organize Exp. 1–6 into three phases [^3] To reproduce Exp. 1–6, please refer to the released artifacts at https://doi.org/10.57760/sciencedb.29613.:
[…]
  • Phase II: Vision intervention for robustness (Secs. V-DV-E). We test whether reducing dominant input–gradient couplings during training improves robustness. Exp. 3 introduces ConjMask, which masks large $|x_{i}p_{i}|$ components on triggered samples, and evaluates robustness under $L_{\infty}$ attacks: PGD-20, APGD-CE, and APGD-DLR. Exp. 4 adds LogitReg to broaden robustness beyond the CE channel, targeting cases where performance is strong under CE-based attacks but weaker under DLR-based attacks (Proposition 3). ConjMask aims to push these samples towards the central “Goldilocks” band in Fig. 1.
  • Phase III: Prefill-only LLM evaluation (Secs. V-FV-G). We test the complementary claim that unusually low prompt-level coupling $c_{\text{prompt}}$ indicates weak conditioning and higher hallucination risk. Exp. 5 evaluates $c_{\text{prompt}}$ as a decoding-free hallucination risk score (AUROC). Exp. 6 applies the same prefill score for prompt selection among paraphrases, measuring whether choosing higher-$c_{\text{prompt}}$ prompts improves judge-preferred selection (Proposition 4). Hallucination-prone prompts reside in the high ambiguity region of Fig. 1 (large boundary thickness, uncontrolled sensitivity).
V-B Exp. 1: CC-Probe Separates Correct and Incorrect Samples During Training
Figure 2: Evolution of the CC-Probe and accuracy during training. We plot evaluation accuracy (red, right axis) and the mean absolute input–gradient cosine $\bar{c}_{\text{img}}$ (left axis; Eq. (33)), computed on the held-out evaluation split every 10 epochs and reported separately for correctly classified versus misclassified samples (green vs. blue). Gradients are taken w.r.t. the standard cross-entropy loss using the ground-truth label. (a) CIFAR-10 (ResNet-18, DenseNet-121, ViT-Tiny, EfficientNet-B0). (b) Tiny-ImageNet-200 (ResNet-50, DenseNet-121, EfficientNet-B4, Swin-Tiny).
For each test input $x$ we compute the per-sample CC-Probe $c_{\text{img}}(x)$ from Eq. (33), where $p(x)=\nabla_{x}\mathcal{L}(f_{\theta}(x),y_{\text{true}})$ is the input gradient of the standard cross-entropy. We then average $c_{\text{img}}(x)$ over the subset of correctly classified samples and over the subset of misclassified samples at each epoch. Since $c_{\text{img}}$ is scale-free, the trends below are not driven by gradient magnitudes; in implementation we stabilize the cosine computation by adding a small value to the input/gradient norm denominators, rather than dropping low-gradient samples.
We tracked $\bar{c}_{\text{img}}$ during training across CNN and Transformer architectures on CIFAR-10 and Tiny-ImageNet-200, recording the statistic separately for correctly and incorrectly classified samples on the evaluation split. As shown in Fig. 2, two consistent phenomena emerged:
  1. (i)
    Correct samples decouple. As training progresses and accuracy improves, $\bar{c}_{\text{img}}$ for correctly classified samples steadily decreases toward zero, indicating that these images tend to decouple with their gradients.
  2. (ii)
    A persistent “boundary-stress”. In contrast, $\bar{c}_{\text{img}}$ for incorrectly classified samples remains elevated, creating a stable gap between the two populations across checkpoints.
A small $\bar{c}_{\text{img}}$ implies weak average conjugate alignment between $x$ and $p(x)=\nabla_{x}\mathcal{L}_{c}(x)$, suggesting that the NUP correlation channel is relatively unsaturated for the corresponding sample set. From the NUP viewpoint, training tends to move correctly classified samples into a high-slack regime (low loss with low residual correlation), whereas misclassified samples remain in a boundary-stress where correlation stays elevated, indicating persistent boundary sensitivity. This pattern supports the redistribution picture in Proposition 1: conjugate alignment concentrates on a small hard subset while most correct samples exhibit low coupling.
V-C Exp. 2: Stepping Along/Against the Gradient Moves the CC-Probe as Predicted
Figure 3: Effect of gradient-aligned perturbations on accuracy and $\bar{c}_{\text{img}}$. We evaluate ResNet-18, ViT-Tiny, and Vim-Tiny on CIFAR-10 and ImageNet-100. Left column: accuracy; right column: $\bar{c}_{\text{img}}$. We compare Clean, +FGSM and –FGSM across perturbation levels $\epsilon$ (see Supplementary Material for exact values).
To test whether $\bar{c}_{\text{img}}$ responds causally to boundary-directed input changes, we evaluate the dedicated Exp. 2 checkpoints (trained with clean data) for ResNet-18, ViT-Tiny, and Vim-Tiny on CIFAR-10 and ImageNet-100. We then perturb each input by a small $L_{\infty}$ step along the sign of the input gradient:
$x^{+}=$ $\mathrm{clip}\!\left(x+\epsilon\,\mathrm{sign}(\nabla_{x}\mathcal{L}_{c}(x))\right),$ (35)
$x^{-}=$ $\mathrm{clip}\!\left(x-\epsilon\,\mathrm{sign}(\nabla_{x}\mathcal{L}_{c}(x))\right),$
where $\mathrm{clip}(\cdot)$ enforces valid pixel ranges. We then recompute $\bar{c}_{\text{img}}$ on the perturbed inputs (i.e., using $(x^{\pm},\,\nabla_{x^{\pm}}\mathcal{L}_{c})$) and evaluate accuracy under the same perturbations.
As illustrated in Fig. 3, we observe a consistent bidirectional response:
  1. (i)
    Gradient-aligned perturbations increase coupling and degrade accuracy. As $\epsilon$ increases, +FGSM typically raises $\bar{c}_{\text{img}}$ while reducing accuracy, consistent with pushing samples toward locally more loss-sensitive regions.
  2. (ii)
    Anti-aligned perturbations reduce coupling (when $\epsilon$ is small). For sufficiently small $\epsilon$, –FGSM often lowers $\bar{c}_{\text{img}}$ and can preserve (or slightly improve) accuracy, reflecting a move toward locally flatter loss regions. For larger $\epsilon$, the perturbation may leave the semantic manifold and accuracy can still drop; our claim here is about the directional effect on the coupling statistic.
In the NUP, proximity to the constraint is governed by the effective conjugate volume $V_{c}(u)=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}$ and the slack ratio $S_{c}(u)=V_{c}(u)/\kappa\geq 1$ (Eq. (28)). Here, the CC-Probe $\bar{c}_{\text{img}}$ is used as a proxy for the input–gradient coupling channel entering $\rho_{c}(u)$ via the loss-induced statistics; it is not itself an exact estimator of $V_{c}(u)$ or $S_{c}(u)$. Accordingly, Fig. 3 is interpreted as an empirical probe of boundary stress: +FGSM increases loss and empirically raises $\bar{c}_{\text{img}}$, which is consistent with stronger correlation pressure and higher boundary stress, whereas sufficiently small –FGSM steps reduce the measured coupling and are consistent with a relaxation toward a less stressed regime.
V-D Exp. 3: ConjMask improves robustness under CE-based first-order attacks
TABLE I: Robust accuracy (%) under $L_{\infty}$ attacks at $\epsilon=8/255$ on CIFAR-10. Attacks: PGD-20, Auto-PGD with CE loss (APGD-CE), Auto-PGD with DLR loss (APGD-DLR). We use APGD-CE/APGD-DLR as objective-specific stress tests (components of AutoAttack), rather than claiming evaluation under the full AutoAttack suite. Timing is reported as provided (seconds $\times 200$), otherwise “–”. Results illustrate NUP-predicted, objective-specific shifts; interventions are demonstrations (not tuned for standardized leaderboard performance). TRADES-PGD is shown only as a reference.
Model Method Clean PGD-20 APGD-CE APGD-DLR Timing (Sec. $\times$ 200)
ResNet-18 Base 91.52 0.01 0.02 0.10
TRADES-PGD 74.57 44.14 43.92 42.27 $\sim$ 100
ConjMask (Exp. 3) 85.27 77.55 77.02 1.28 $\sim$ 20
ConjMask+LogitReg (Exp. 4) 87.64 62.71 36.92 43.29 $\sim$ 30
ViT-Tiny Base 84.10 0.01 0.01 0.02
TRADES-PGD 71.95 32.95 32.46 35.51 $\sim$ 120
ConjMask (Exp. 3) 71.49 39.19 29.49 13.96 $\sim$ 45
ConjMask+LogitReg (Exp. 4) 71.24 43.15 14.59 15.91 $\sim$ 60
EfficientNet-B0 Base 88.14 0.24 0.17 0.00
TRADES-PGD 56.88 30.65 30.48 28.05 $\sim$ 260
ConjMask (Exp. 3) 83.91 69.10 68.28 0.20 $\sim$ 50
ConjMask+LogitReg (Exp. 4) 81.18 48.48 21.89 28.36 $\sim$ 65
Following Proposition 3, we study whether a training-time suppression of highly conjugated input components can improve adversarial robustness without adversarial training. Here, ConjMask is an objective-specific test of NUP, rather than a solution of comprehensive adversarial robustness.
We start from a separately trained clean CIFAR-10 checkpoint for the same architecture. For each mini-batch, we first compute a forward pass and mark samples as triggered if either the current prediction is incorrect, or the prediction confidence is below a fixed threshold (default $0.2$). For each triggered image $X$ (flattened as $x$) we compute the input gradient $P=\nabla_{X}\mathcal{L}$ (flattened as $p$). In the released implementation, this probing step intentionally switches the model to eval mode; moreover, this mode is kept for the subsequent forward passes of the same optimization step. Since the goal of ConjMask is to reduce the CC-Probe $c_{\text{img}}=|\cos(x,p)|$ (Eq. (33)), we focus on the components that contribute most to this coupling statistic.
To do so, we build a per-pixel importance score by taking the elementwise product between unit-normalized image and gradient (done per channel in the implementation):
$s_{c,j}=\big|\tilde{x}_{c,j}\,\tilde{p}_{c,j}\big|,\qquad\tilde{x}_{c}=\frac{x_{c}}{\|x_{c}\|_{2}+\epsilon},\ \tilde{p}_{c}=\frac{p_{c}}{\|p_{c}\|_{2}+\epsilon}.$ (36)
We then select the top-$k$ pixels per channel according to $s_{c,j}$ and first form a binary mask. In the released implementation, this mask is subsequently smoothed with a Gaussian kernel to obtain a soft mask $M\in[0,1]^{C\times H\times W}$, and the triggered input is replaced by soft stochastic interpolation between the original image and a per-channel Gaussian replacement sampled from CIFAR-10 statistics:
$X_{\mathrm{mask}}=X\odot(1-M)+R\odot M.$
Thus, ConjMask is a conjugation-guided soft stochastic replacement.
We report robust accuracy under $L_{\infty}$ attacks at $\epsilon=8/255$, using PGD-20 and Auto-PGD variants (CE and DLR losses), together with standard clean accuracy in Table I. The following observations can be made:
  1. (i)
    ConjMask induces a large regime shift under CE-based first-order attacks. On CIFAR-10 at $\epsilon=8/255$ under PGD-20 and APGD-CE, clean baselines largely collapse, while ConjMask (Exp. 3) substantially improves robustness on CNNs and without any test-time masking. Importantly, the robustness profile is objective-dependent: the same models can remain vulnerable under APGD-DLR, which we interpret as evidence that ConjMask primarily manipulates the CE-gradient coupling channel predicted by the NUP, rather than constituting a universally strong defense across objectives.
  2. (ii)
    CE robustness can remain objective-dependent. The same models that are strong on PGD/APGD-CE can still degrade sharply under APGD-DLR: for ConjMask (Exp. 3), APGD-DLR robust accuracy is $1.28\%$ on ResNet-18, $13.96\%$ on ViT-Tiny, and $0.20\%$ on EfficientNet-B0 (Table I). This “strong on CE, weaker on DLR” pattern highlights objective-specific geometry channels, which is consistent with the NUP view that manipulating one coupling channel may not cover all attack objectives.
  3. (iii)
    ViT benefits less from ConjMask than CNNs. For ViT, Exp. 3 is stronger than the clean baseline under CE-based attacks, but it does not surpass TRADES-PGD. We attribute this gap to architectural differences: compared with CNNs, ViTs may exhibit weaker feature compression and preserve more non-pixel-level information pathways, which can reduce the effectiveness of a pixel-level masking intervention.
Importantly, all masking is used only during fine-tuning; at test time, the deployed model always receives the original clean input, with no masking, no stochastic preprocessing, and no gradient-obfuscation mechanism.
V-E Exp. 4: LogitReg complements ConjMask beyond the gradient channel
Exp. 3 demonstrates that conjugation-guided masking can produce strong resistance to standard gradient attacks, but its near-zero APGD-DLR accuracy indicates a remaining vulnerability that is not explained by the CE-gradient channel alone.
In Exp. 4, we keep the same ConjMask masking procedure as in Exp. 3 and add an output-side regularization term, denoted by LogitReg, together with an optional auxiliary-view consistency term. The purpose is to complement the CE-gradient coupling channel targeted by ConjMask with score-space stabilization, so that robustness is not restricted to CE-like attack directions.
Concretely, for Exp. 4 the objective is
$\mathcal{L}_{\text{Exp4}}=\mathcal{L}_{\text{mask}}+\mathcal{L}_{\text{logit}}+\mathcal{L}_{\text{cons}},$ (37)
where $\mathcal{L}_{\text{mask}}$ is the same masked-input cross-entropy term as in Exp. 3:
$\mathcal{L}_{\text{mask}}=\mathrm{CE}\!\big(f_{\theta}(X_{\mathrm{mask}}),y\big).$ (38)
The released script uses a shared logit-side stabilizer across ResNet-18, ViT-Tiny, and EfficientNet-B0, namely a centered-logit variance penalty. Let
$\bar{z}=z-\frac{1}{K}\sum_{j=1}^{K}z_{j},\qquad\mathrm{VarC}(z)=\frac{1}{K}\sum_{j=1}^{K}\bar{z}_{j}^{2}.$ (39)
Then
$\mathcal{L}_{\text{logit}}=\alpha_{\mathrm{logit}}\cdot\frac{1}{2}\Big(\mathrm{VarC}(f_{\theta}(X))+\mathrm{VarC}(f_{\theta}(X_{\mathrm{mask}}))\Big),$ (40)
where $\alpha_{\mathrm{logit}}=0$ at epoch 0, and $\alpha_{\mathrm{logit}}=\texttt{penalty\_weight}$ thereafter.
When the consistency branch is activated, the auxiliary view is a per-sample mean-baseline image:
$X_{\mathrm{aux}}=\bar{X}\mathbf{1},\qquad\bar{X}=\frac{1}{CHW}\sum_{c,h,w}X_{c,h,w}.$ (41)
The consistency term is
$\mathcal{L}_{\text{cons}}=0.8\,\mathrm{KL}\!\Big(\log\mathrm{softmax}\!\big(\tfrac{f_{\theta}(X_{\mathrm{mask}})}{T}\big)\,\big\|\,\mathrm{softmax}\!\big(\tfrac{f_{\theta}(X_{\mathrm{aux}})}{T}\big)\Big),$ (42)
with $T=1$. In the released implementation, this branch is enabled for epochs greater than 5.
We evaluate under the same $L_{\infty}$ setting as Exp. 3 ($\epsilon=8/255$) using PGD-20, APGD-CE, and APGD-DLR. The following observations can be made:
  1. (i)
    Lightweight stabilizers restore DLR robustness with practical efficiency. Adding shared logit regularization and shared auxiliary-view consistency on top of the same ConjMask procedure substantially improves APGD-DLR robustness where Exp. 3 remains weak. On ResNet-18, APGD-DLR increases from $1.28\%$ (Exp. 3) to $43.29\%$ (Exp. 4), close to TRADES-PGD’s $42.27\%$, while maintaining high clean accuracy ($87.64\%$) and requiring no adversarial training, hence lower practical cost. The resulting robustness profile is more balanced than Exp. 3, with non-trivial performance retained under PGD-20 and APGD-CE ($62.71\%$ and $36.92\%$, respectively).
  2. (ii)
    The gain is architecture-dependent in outcome, reflecting CE–DLR direction mismatch with partial overlap. For ViT-Tiny and EfficientNet-B0, Exp. 4 trades part of the PGD/APGD-CE robustness of Exp. 3 for improved APGD-DLR robustness (ViT-Tiny: $13.96\%\to 15.91\%$; EfficientNet-B0: $0.20\%\to 28.36\%$), but the improvement remains less pronounced than on ResNet-18 and is accompanied by a noticeable drop in PGD-20/APGD-CE.
Overall, the shared LogitReg augmentation can deliver non-trivial robustness improvements with much less training time than adversarial training.
V-F Exp. 5: Prefill-only hallucination risk prediction on Benchmark-500
Here, we use “hallucination” in an operational sense: a response is labeled as hallucinated if it contains confident but unsupported/incorrect reasoning or claims, as judged by external verifiers. In math reasoning, this includes fabricated intermediate steps, invalid transformations, or incorrect final answers presented with high confidence. We emphasize that this is a task-specific reliability label used for evaluating the proposed prefill-only risk signal, rather than a universal definition of hallucination.
We intentionally restrict to a strict prefill-only, single-backward setting: all risk scores are computed before generating any answer tokens. Under this constraint, we do not aim to compete with post-hoc hallucination detectors that rely on sampling multiple generations, retrieval, tool use, or LLM-as-a-judge at test time; instead we test whether a pre-hoc geometric signal exists as predicted by the NUP.
We test Proposition 4 by testing whether low prompt-level conjugate correlation signals hallucination risk in a trained LLM. For each of 500 math problems, we compute the prefill-stage probe on the prompt embeddings: we obtain the embedding tensor $x$ and its input gradient $p=\nabla_{x}\mathcal{L}(x)$ from a single backward pass, and form $c_{\text{prompt}}$ (Eq. (34)), which is computed before any answer tokens are generated.
We then generate a single model response only for evaluation, and use a panel of external LLM judges to assign a hallucination vote probability $P_{\text{vote}}\in[0,1]$ for that response.
Risk scores
We evaluated four sample-level probes as hallucination risk scores (higher = more likely to hallucinate):
  • Risk-Cos (ours): $c_{\text{prompt}}$;
  • Risk-Entropy: mean predictive entropy;
  • Risk-Loss: mean negative log-likelihood (shifted NLL as defined in Sec. IV-B);
  • Risk-Margin: negative mean logit margin.
Note that Risk-Loss here reflects prompt-level self-supervised NLL, not factual correctness of the final answer; thus it can behave very differently from a hallucination indicator.
Evaluation
We report AUROC with bootstrap resampling ($N_{\text{boot}}=2000$) for distinguishing strict-consensus hallucinations ($P_{\text{vote}}=1.0$) from strict-consensus clean responses ($P_{\text{vote}}=0.0$), and also provide a relaxed-consensus analysis ($P_{\text{vote}}\geq 0.8$) as a robustness check. Results are summarized in Table II. We can observe that:
  1. (i)
    Standard uncertainty metrics fail. Risk-Entropy is near chance, while Risk-Loss is often anti-correlated with hallucination labels, highlighting that hallucinations can be fluent and low-entropy rather than “uncertain”.
  2. (ii)
    Risk-Cos is effective as a prefill-stage geometric signal. Risk-Cos achieves AUROC $\approx 0.69$ under both strict and relaxed consensus settings. This supports the slack-regime interpretation of Proposition 4: hallucinations correspond to anomalously low prompt–gradient alignment, indicating weak conditioning and a larger feasible continuation space prior to decoding.
TABLE II: AUROC of prefill-stage hallucination risk signals on Benchmark-500. All metrics are computed without generating any answer tokens (prefill only). Risk-Cos (ours) and Risk-Margin are informative, while prompt-level entropy and NLL are weak under this constraint.
Metric AUROC Mean 95% CI Interpretation
Strict Consensus ($P_{\text{vote}}=1.0$ vs. Clean)
Risk-Cos (Ours) 0.6939 [0.58, 0.79] Informative
Risk-Margin 0.6850 [0.57, 0.79] Informative
Risk-Entropy 0.5323 [0.44, 0.63] Near chance
Risk-Loss 0.3508 [0.25, 0.46] Anti-correlated
Relaxed Consensus ($P_{\text{vote}}\geq 0.8$)
Risk-Cos (Ours) 0.6957 [0.61, 0.78] Robust
Risk-Margin 0.6730 [0.59, 0.75] Robust
Risk-Entropy 0.5020 [0.42, 0.59] Near chance
Risk-Loss 0.3260 [0.25, 0.41] Anti-correlated
Mechanistic interpretation under Proposition 4
Under the prefill-only constraint, both Risk-Entropy and Risk-Loss are functions of the model’s next-token distribution over the prompt itself; thus they primarily quantify the predictability of the prompt wording under the model rather than the correctness of the (future) solution. Their weak/negative correlation indicates that “easy-to-model” prompts (low $\mathrm{NLL}_{\text{prompt}}$ or low entropy) can still admit confident but incorrect continuations, i.e., prompt predictability does not control downstream factuality.
In contrast, Risk-Cos targets the conjugate coupling channel posited in Proposition 4. Since $c_{\text{prompt}}$ serves as a computable surrogate for the input–gradient covariance channel that enters $\rho_{c}(u)$ in the NUP, a larger coupling magnitude (larger $|\rho_{c}(u)|$) compresses the effective conjugate volume through the factor $\sqrt{1-\rho_{c}(u)^{2}}$, yielding smaller slack $S_{c}(u)$ (stronger conditioning), whereas a smaller coupling corresponds to a higher-slack regime (weaker conditioning). Thus, low $c_{\text{prompt}}$ (high Risk-Cos) indicates that the prompt content is weakly coupled to the loss-sensitive directions, consistent with Proposition 4’s under-conditioning hypothesis: a high-slack state admits a larger set of feasible continuations prior to decoding, increasing the likelihood of prior-driven drift that manifests as hallucination.
V-G Exp. 6: Prefill-stage Prompt Selection
We evaluate whether a conjugate-correlation score can be used as a prefill-only quality filter to select a prompt before generating any answer tokens.
Exp. Setup
We use Perturbation-100, consisting of 100 problems with 5 semantically equivalent (paraphrased) prompt variants per problem. For each problem, we compute a prefill risk score for all five candidate prompts without decoding any answer tokens. The selection rule itself is therefore prefill-only: for each metric, the prompt with the lowest predicted risk is designated as the selected variant. For offline evaluation, we then generate deterministic answers for all five prompt variants and score them with the same judge panel as in Exp. 5. The resulting judged answers are used only to identify the judge-best set for each problem (allowing ties among highest-scoring variants) and to compute the Top-1 / Mean-Regret metrics reported below. As a sanity check on the reliability of the judge panel, we also report inter-judge agreement via Fleiss’ Kappa.
Prefill-only risk scores.
Each candidate prompt is assigned one scalar risk value (higher $\rightarrow$ worse) computed before decoding: Risk-Cos (ours), Risk-Margin, Risk-Entropy, and Risk-Loss (NLL) (definitions follow Sec. IV-B). All reported selection decisions in this experiment use only these prefill scores.
Metrics
We focus on two selection-oriented metrics that directly reflect the practical utility of a prefill-only prompt filter:
  • Top-1 Hit Rate ($\uparrow$): the fraction of problems for which the selected prompt belongs to the judge-best set.
  • Mean Regret ($\downarrow$): the average quality gap between the selected prompt and the best judge score attained within the same five-prompt cluster.
In addition, we report Fleiss’ Kappa on the judges’ binary hallucination annotations as a sanity check on inter-judge consistency.
Prefill Risk Score Top-1 Hit Rate $\uparrow$ Mean Regret $\downarrow$
Risk-Cos (ours) 0.83 0.433
Risk-Margin 0.81 0.591
Risk-Loss (NLL) 0.78 0.615
Risk-Entropy 0.77 0.636
TABLE III: Prefill-only prompt selection results on Perturbation-100. For each five-prompt cluster, the prompt with the lowest predicted prefill risk is selected before decoding. Risk-Cos achieves the highest Top-1 Hit Rate and the lowest Mean Regret among all compared methods.
Results
Table III shows that Risk-Cos achieves the highest Top-1 Hit Rate (0.83) and the lowest Mean Regret (0.433) among all compared prefill-only risk scores. This indicates that, under the prefill-only constraint, Risk-Cos is the strongest criterion on the two primary selection metrics reported in the main text.
Risk-Margin attains a similarly high Top-1 Hit Rate (0.81), but its Mean Regret is larger (0.591). This gap suggests that although Margin often selects a competitive prompt, its failures are more costly when it misses the judge-best variant. By comparison, Risk-Loss (NLL) and Risk-Entropy perform worse on these two primary metrics, with lower hit rates (0.78 and 0.77, respectively) and higher regret (0.615 and 0.636, respectively).
Judge reliability
The judge panel exhibits strong agreement on the binary hallucination annotation, with Fleiss’ Kappa = 0.857. This serves as a sanity check that the judges are broadly consistent on that binary label and supports the stability of the offline evaluation pipeline. We emphasize, however, that the prompt-ranking score used in Exp. 6 is an aggregate judge score that also depends on additional structured judge fields (see Supplementary Material), so Fleiss’ Kappa should not be interpreted as a direct reliability coefficient for that full aggregate score.
Overall, the results of Exp. 6 suggest that Risk-Cos can serve as a useful, decoding-free criterion for selecting among semantically equivalent prompt formulations. More importantly, this finding provides direct empirical support for the NUP account developed in this paper: within the prefill stage, stronger prompt–gradient coupling is associated with better-constrained continuations, whereas weaker coupling reflects the under-conditioning regime predicted by Proposition 4.
VI Discussion
The primary theoretical value of the NUP is that it provides a fundamental boundary on what a learned system can simultaneously achieve when its behavior is examined through the joint geometry of inputs and loss sensitivities. In the Robertson–Schrödinger form (Theorem III.1), the pair $(\hat{x}_{u},\hat{p}_{u})$ obeys an irreducible constraint: the boundary-emphasized phase-plane area cannot collapse below a constant, equivalently $\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\geq\tfrac{1}{2}$ (Lemma III.1). This plays a role analogous to a Carnot-type limit: regardless of architecture or training details, there exists a nonzero “minimum uncertainty budget” that prevents simultaneously making boundary ambiguity arbitrarily small and boundary sensitivity uniformly low. In this sense, the NUP is not merely descriptive; it states an intrinsic incompatibility that any robust-and-accurate (or faithful-and-confident) system must navigate.
VI-A A conjugate limit underlying boundary behaviors
Interpreting the operator moments as second-order geometry of the loss-weighted $(x_{u},p_{u})$ cloud (Sec. III-D) turns the inequality into a geometric statement: the effective area $\sqrt{\det\Sigma(u)}=\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}$ is lower-bounded (Eq. (26)). When the system operates near this limit, improvements in one axis necessarily incur costs in the conjugate axis. This offers a principled explanation for why purely optimizing one desideratum (e.g., squeezing boundary errors to raise accuracy) can systematically amplify another failure mode (e.g., sensitivity exploited by adversarial perturbations).
VI-B Unified interpretation across vision and language
Within this limit-based view, adversarial fragility and hallucination become two opposite ways of mismanaging the same uncertainty budget. Table IV summarizes how NUP manifests differently across modalities.
TABLE IV: Two regimes from the NUP. The NUP constrains an effective conjugate volume $V=\Delta x\,\Delta p\,\sqrt{1-\rho^{2}}\geq\kappa$, where $\rho$ is the operator correlation between $\hat{x}_{u}$ and $\hat{p}_{u}$ in the RS inequality. Empirically, we diagnose the input–gradient coupling channel entering $\rho$ using the $C$-Probe: in vision we use $|\cos(x,p)|$ (inputs are normalized), while in LLM prompting we use the centered $|\cos(\bar{x},\bar{p})|$.
Saturation / Boundary-stress (Vision) Slack / Under-conditioning (LLM)
NUP mechanism $|\rho|\uparrow\Rightarrow\sqrt{1-\rho^{2}}\downarrow\Rightarrow V\downarrow\$(tight, near-bound) $|\rho|\downarrow\ (\to 0)\Rightarrow\sqrt{1-\rho^{2}}\uparrow\Rightarrow V\uparrow\$(loose, high slack)
Probe signal high $|\cos(x,p)|$ low $|\cos(\bar{x},\bar{p})|$
Observed symptom misclassification & adversarial fragility higher hallucination risk
Actionable interventions ConjMask + LogitReg Prefill Risk Scoring + Prompt Screening/Selection
In discriminative vision, training can push hard samples toward a boundary-stress configuration: the ambiguity thickness becomes small while sensitivity dispersion becomes large, with strong input–gradient coupling concentrating on a stressed tail (Exp. 1–2; Proposition 12). In LLM prompting under a prefill-only constraint, the dominant pathology is instead under-conditioning: unusually weak prompt–gradient coupling indicates that the prompt does not effectively constrain loss-sensitive directions, leaving high slack and enabling prior-driven drift (Exp. 5–6; Proposition 4). The same coupling channel therefore provides a common language for boundary anomalies across perception and generation.
VI-C From a bound to practical observables and control channels
A second theoretical contribution is the bridge from the operator inequality to computable quantities. Under the loss-induced state, the symmetrized operator covariance reduces to a scalar covariance of real gradients (Theorem III.2), and random directional probing links projected correlation to cosine (Lemma III.2). This justifies the CC-Probe as a single-backward proxy for the coupling channel that modulates proximity to the uncertainty limit. Once coupling is made observable, it also becomes controllable: ConjMask directly suppresses dominant per-component couplings $|x_{i}p_{i}|$ and yields strong robustness under PGD-20 and APGD-CE without adversarial training (Exp. 3; Proposition 3). The CE–DLR gap further indicates that robustness is multi-channel; LogitReg complements ConjMask by stabilizing score-space geometry and broadening robustness under APGD-DLR (Exp. 4).
VI-D Limitations and future directions
Our analysis is most informative when there exists a nontrivial boundary-relevant subset that the loss-phase construction can emphasize. In the language setting, the usefulness of CC-Probe depends on a nontrivial match between benchmark difficulty and model capability: if the benchmark is too easy hallucinations are too rare, if it is too hard failures become nearly ubiquitous, and the probe is therefore most informative on prompts near the model’s effective uncertainty boundary.
In near “pure fitting” regimes—e.g., when the loss is essentially small on the relevant support, or the input–output mapping behaves close to bijective with little effective compression [40]—the weighting $w_{c}(x)\propto\mathcal{L}_{c}(x)^{2}$ provides little contrast and distinct loss-induced states $\psi_{c}$ are not meaningfully separated; consequently, the notion of a stressed boundary layer becomes ill-posed and the CC-Probe inclines to lose discriminative power. In addition, the probe is loss-dependent (CE in vision; shifted prompt NLL in language), and Exp. 3 indicates robustness can differ across attack losses (CE vs. DLR), motivating multi-loss evaluation (e.g., APGD-CE and APGD-DLR) and defenses that cover multiple geometry channels. Finally, CC-Probe is a proxy rather than a direct estimator of $\rho_{c}(u)$ or slack $S_{c}(u)$, and prefill-only LLM use still requires one backward pass on prompt embeddings. Future work includes designing alternative state constructions when $\mathcal{L}_{c}$ is near-degenerate, developing stable estimators of direction-dependent slack, and extending coupling-based monitoring/control to multimodal and preference-optimized settings.
Crucially, the NUP framework also serves as a predictive warning for future LLM alignment. Currently, LLM hallucinations are predominantly driven by a high-slack (under-conditioned) regime, as detected by the CC-Probe. However, if future training paradigms aggressively reduce this slack to suppress hallucinations—forcing prompts into tight, deterministic couplings with specific continuations—the NUP inequality $\Delta\hat{m}_{u}^{\star}\,\Delta\hat{p}_{u}\geq\tfrac{1}{2}$ dictates that the system could inevitably shift toward the opposite extreme. In doing so, we risk inducing the same boundary-stress regime currently plaguing vision models, which would manifest in LLMs as extreme adversarial fragility (e.g., imperceptible prompt perturbations or synonymous substitutions triggering catastrophic output shifts). Ultimately, the NUP demonstrates that we cannot entirely eliminate boundary uncertainty; we can only trade one form of fragility for another. Sustainable model design might therefore navigate this fundamental trade-off, seeking an intermediate “Goldilocks” zone rather than minimizing a single failure axis.
VII Conclusion
In this study we have introduced the Neural Uncertainty Principle (NUP), a fundamental geometric constraint showing that input projections and their loss gradients are conjugate observables under a loss-induced state. NUP unifies adversarial vulnerability in vision and hallucination in LLMs as opposite extremes on the same conjugate plane—saturation versus slack. Based on this principle, we derive a practical single-backward probing method, CC-Probe, which enables the diagnosis of cross-modal boundary anomalies and supports targeted interventions: ConjMask and LogitReg for enhancing robustness in vision tasks, and prefill-stage risk scoring for hallucination detection and prompt selection in language models. By reframing two isolated reliability challenges as a single conjugate trade-off, the NUP provides a unified perspective for advancing robust and faithful AI.
Reproducibility and Artifact Release
Training/eval scripts, seeds, logs, and prompt lists are released at https://doi.org/10.57760/sciencedb.29613. Hardware budgets (GPU-days), wall-clock, memory footprints, exact hyperparameters, and attack settings are included in the Supplementary Material.
Acknowledgments
During manuscript preparation, the authors used ChatGPT (OpenAI) for language editing, cross-section consistency checking, and suggestions aimed at improving the clarity and presentation of mathematical exposition. All technical content, including the underlying proofs, derivations, experimental design, implementation, and conclusions, was developed and verified by the authors, who take full responsibility for the final manuscript.